# AI-enhanced Collective Intelligence: The State of the Art and Prospects

### v2

@@ 2.3.1. Surface-level diversity
The evidence concerning the effect of gender diversity on team performance is equivocal and contingent upon various contextual factors (Bear and Woolley, 2011). Past research found that gender-diverse teams outperform gender-homogeneous teams when perceived time pressure is low (Kearney et al., 2022). One study showed gender diversity was negatively associated with performance, but only in large groups (Wegge and Schmidt, 2009). Previous evidence strongly suggests that team collaboration is greatly improved by the presence of females in the group (Bear and Woolley, 2011; Woolley et al., 2010). Gender-diverse teams produce more novel and higher-impact scientific ideas (Yang et al., 2022). Other studies show that gender composition seems to matter for team performance. However, when controlling for the individual abilities of team members, the relation between gender composition and team performance vanishes (Städter et al., 2022). One study points out that the ”romance” of working together can benefit group performance (Zhang and Hou, 2012). Gender-diverse groups perform better than homogeneous groups by decreasing relationship and task conflict (Zhang and Hou, 2012).
@@
Age diversity may potentially enhance group performance, but the benefits may depend on the context. Several studies have found that age diversity in teams positively correlated with team performance when completing complex tasks (Wegge et al., 2012).
Past research points out the positive effects were only seen under conditions of positive team climate and low age discrimination (Wegge et al., 2012). Another study shows teams with more age diversity reported more age discrimination (Kunze et al., 2013), associated with lower commitment and worse performance.
@@
There are inconsistent results in the case of ethnic diversity. One study (McLeod et al., 1996) found racially diverse groups produced significantly more feasible and effective ideas than homogeneous groups. Another study found racial similarity in groups associated with higher self-rated productivity and commitment (Riordan and Shore, 1997). Under the right conditions, teams may benefit from diversity in ethnicity and nationality (Nielsen et al., 2017).
AI exhibits surface-level diversity through different anthropomorphic features in robots, such as voice, avatars, and human-like characteristics. Research has indicated that the gendering of AI can affect perceptions and trust levels in users. For instance, studies have found that users may exhibit more trust in AI personal assistants whose voice gender matches their own (Lee et al., 2021). Assigning a female gender to AI bots enhances their perceived humanness and acceptance (Borau et al., 2021). Previous studies have indicated a perception bias in AI agents based on gender: male AI agents tend to be viewed as more competent, whereas female AI agents are often perceived as warmer (Ahn et al., 2022). The perceived gender of the machine can make the social dynamics in hybrid teams even more complicated. However, these gendered characteristics in AI also raise concerns about reinforcing stereotypes and biases (Chin and Robison, 2020). The influence of AI’s gendered traits on human interactions is an ongoing area of research, with implications for how AI is designed and utilized.
2.3.2. Deep-level diversity
@@
Some work proposes that investigations into the relationship between personality and work-related behaviors should expand beyond the linearity assumptions (Curşeu et al., 2019), showing that extraversion, agreeableness, and conscientiousness have inverted U-shaped relationships with peer-rated contributions to teamwork (Curşeu et al., 2019). Similar work demonstrates a curvilinear relationship between a team’s average proactive personality and performance (Zhang et al., 2021a). This body of work also identifies the moderating effects of personality dispersion, team potency, and cohesion on this relationship (Zhang et al., 2021a). Previous research has found that a team’s collective openness to experience and emotional stability moderates how task conflict affects team performance (Bradley et al., 2013). Specifically, in teams with high openness or emotional stability, task conflict positively influences performance (Bradley et al., 2013).
@@
Team cognitive style diversity refers to the variation in team members’ ways of encoding, organizing, and processing information (Aggarwal et al., 2019). Research indicates its significance in fostering innovation within teams (Aggarwal and Woolley, 2019). A certain level of cognitive diversity contributes to collective intelligence by providing diverse cognitive inputs and viewpoints necessary for task work (Aggarwal and Woolley, 2019). However, excessive diversity may lead to high coordination costs (Straub et al., 2023), as team members with differing perspectives struggle to understand each other (Cronin and Weingart, 2007; Mello and Rentsch, 2015). Studies suggest that the relationship between cognitive style diversity and collective intelligence is curvilinear, forming an inverted U-shape (Aggarwal et al., 2019). Additionally, the beneficial effects of cognitive diversity might be mediated through factors such as task reflexivity and relationship conflict (Chen et al., 2019).
@@
Values are internalized beliefs that can guide behavior and enhance motivation (Cennamo and Gardner, 2008). Past research found higher diversity in values was associated with more conflict (Liang et al., 2012) in relationships, tasks, and processes. Overall, existing research consistently indicates that value congruence within teams positively influences performance, satisfaction, and conflict (Jansen and Searle, 2021) and moderates the effects of informational diversity (Jehn et al., 1999).
@@
While AI systems can display various behaviors or responses based on their programming and training, this diversity is a product of different algorithms, models, and data inputs. Unlike humans, AI does not have personal experiences, beliefs, or innate personality traits that contribute to deep-level diversity. In contrast, AI might influence deep-level diversity in human teams, primarily by shaping social interactions and decision-making processes. The optimal composition of teams remains uncertain (Bell, 2007). A previous review underscores the need for in-depth research into both surface and deep-level diversity within teams (Jansen and Searle, 2021). This research should explore how these aspects of diversity influence various outcomes, examining their effects over time and across different levels in a broad array of contexts (Jansen and Searle, 2021).
2.4. Interactions
@@
Interaction distinguishes a group from a mere collection of individuals (Janssens et al., 2022): one person’s behavior forms the basis for the responses of others (Driskell and Salas, 1992). Therefore, the interaction between two members establishes a link in the interaction network. In terms of the group’s interaction network structure, past research suggests that the small-worldness of a collaboration network (small network diameter and high clustering) improves performance up to a certain point, after which the effect will reverse (Uzzi and Spiro, 2005). An ”inverted U” relationship is found between connectedness and performance (Lazer and Friedman, 2007), suggesting there is an optimum number of connections for any given size of the collaboration network. A modular structure (a large number of connections within small groups that are loosely connected) can increase efficiency (Navajas et al., 2018). A study on forbidden triads among jazz groups shows that heavily-sided open triangles are associated with lower success (measured by the number of group releases) (Vedres, 2017).
@@
AI can contribute to human groups in various ways by augmenting existing human skills or complementing capabilities that humans do not possess. AI has attributes beyond humans, such as more extended memory, higher computational speed, and a more vital ability to work with large, diverse datasets. AI’s augmentation of human cognition enables teams to navigate complex situations effectively (Woolley et al., 2023b). Additionally, AI may also help reduce the harm of implicit bias in humans (Lin et al., 2021) and shape better decisions. Taking advantage of the differences between humans and AI can contribute to the performance of hybrid human-AI groups. In the following sections, we discuss the rules and incentives that govern the interaction processes and the structure and dynamics of the group’s interaction network.
2.4.1. Rules and incentives
@@
Decision rules function as prescribed norms that direct interactions and play a crucial role in shaping the communication and information integration within a group (Meslec et al., 2014). Previous research has focused on disentangling decision rules guiding the team’s interaction, ultimately fostering synergy. Some literature emphasizes that the decision rules themselves are intelligent (Wolf et al., 2015). In addition to the rules guiding behavior, the group members interact with each other under specific incentive schemes. The intrinsic and extrinsic incentives are the driving forces (Giacomelli, [n. d.]) for the agents to behave in a direction such that the social network becomes dynamic and moves towards a collective goal. Some research proposes that an incentive scheme rewarding accurate predictions by a minority can foster optimal diversity and collective predictive accuracy (Mann and Helbing, 2017).
@@
AI can augment human cognition (Regens, 2019) to help teams adapt to complexity. While AI can simulate certain human-like incentives through programming and learning algorithms (Park et al., 2023), fully capturing the depth and nuance of human behavior, such as irrationality or altruism, is challenging. A recent study applying a Turing test to compare AI chatbots and humans found that AI chatbots exhibited more altruistic and cooperative behavior than their human counterparts in behavioral games (Mei et al., 2024). Machines can be programmed to mimic human behaviors, but whether they genuinely ”exhibit” them like humans is debatable. In the context of collective intelligence, integrating such human-like behaviors in AI may enhance collaboration and decision-making within diverse teams. However, the extent to which AI can be truly self-aware and authentically replicate complex human traits remains a subject of ongoing research and philosophical debate.
2.4.2. Group structure and dynamics
@@
Much research on collective intelligence has been dedicated to finding a group’s optimum size and structure. Several studies have reported that performance improves with group size through enhanced diversity (Malone and Bernstein, 2022), whereas others suggest the opposite: small groups are more efficient (Galesic et al., 2018). Another study suggests that large teams develop and small teams innovate (Wu et al., 2019). Larger groups can better utilize diverse perspectives and knowledge to solve complex problems but also tend to experience more coordination problems and communication difficulties (Woolley et al., 2010). The consideration of optimal group size can be related to the dynamics of the group, whether it is static, growing, or diminishing.
In human-AI hybrid teams, the concept of group size becomes nuanced with AI integration, especially when AI entities such as Large Language Models (LLMs) are involved. The countability of AI agents, particularly if multiple agents are derived from a single LLM, raises questions about their distinctiveness and individuality in a team context. Unlike humans, AI agents can simultaneously process multiple tasks, blurring the traditional boundaries of group size. Applying past research on human group dynamics to these hybrid systems requires rethinking the notions of individual contribution, team cohesion, and communication. It is important to consider how AI’s unique capabilities and scale impact these dynamics and whether multiple AI agents from a single model represent distinct entities or a collective intelligence resource.
@@
research shows that network structure, specifically hierarchy, and link intensity, can affect information distribution (Tang and Liao, 2023), influencing group performance. Another study observed that the presence of structural holes (Burt, 2018) in leaders’ networks and the adoption of core-periphery and hierarchical structures in groups correlated negatively with their performance (Cummings and Cross, 2003). Incorporating the temporal dimension, adaptive social networks can significantly impact collective intelligence by allowing the network’s structure to evolve based on feedback from its members (Almaatouq et al., 2020).
Team size moderates the relationship between network structure and performance (Yuan and Van Knippenberg, 2022). Studies have found central members can coordinate with other team members more easily in smaller teams (Balkundi and Harrison, 2006), whereas in larger teams, communication challenges may arise (Hollenbeck et al., 2002). A recent study shows a more central connected leader in advice-giving networks has a more positive impact on the performance of larger teams, but this effect is reversed in smaller teams (Yuan and Van Knippenberg, 2022).
Determining the optimal size, structure, and human-AI ratio for hybrid teams is a complex task requiring further study, particularly concerning task-specific requirements. Moreover, incorporating AI into human teams can reshape group hierarchies, norms, and rules, thereby affecting collective intelligence.
@@
Teams often outperform individuals primarily due to explicit communication and feedback within the team (Larson Jr, 2013). Understanding the communication network in crowds is crucial for effectively designing and managing crowdsourcing tasks (Yin et al., 2016). Various studies highlight that centralized and decentralized communication patterns can effectively promote team performance, contingent on the nature of the task and the team’s composition (Cohen et al., 1962; Rosen, 2008). In particular, teams handling complex tasks tend to be more productive with decentralized communication networks (Shaw, 1964). Conversely, the centralization of communication around socially dominant or less reflective individuals can negatively impact the utilization of expertise and overall team performance (Sherf et al., 2018). Communication patterns, such as equal participation and turn-taking of team members, can positively affect collective intelligence (Woolley et al., 2010).
@@
The integration of AI in human groups can influence the conversation dynamics in different ways, depending on the role of AI. AI chat interventions can improve online political conversations at scale (Argyle et al., 2023), reducing the chance of conflicts. While AI machines might struggle with capturing the subtle and ineffable social expressions that make up the dynamics of human groups, this can be advantageous. In certain contexts, AI can be social catalysts (Rahwan et al., 2020) to promote communication where human capacity is limited. Striking the right balance in AI’s role is key to maintaining the natural dynamics of human interaction.
Cognition processes
@@
Cognitive processes are foundations for the emergence of collective intelligence (Woolley and Gupta, 2023), encompassing knowledge acquisition, information processing, problem-solving, decision-making, language, perception, memory, attention, and reasoning. Human cognition processes evolve and interact with each other during various tasks and experiences. Humans can represent the mental states of others, including desires, beliefs, and intentions (Rabinowitz et al., 2018), known as the Theory of Mind (ToM). Past research, inspired by ToM, showed that the ability of team members’ to assess the mental state of others is positively associated with team performance both in face-to-face settings (Woolley et al., 2010) and online (Engel et al., 2014). Regarding collective understanding within a group, the Shared Mental Models (SMM) theory is helpful for understanding, predicting, and improving performance in human teams (Andrews et al., 2023).
@@
Applying theories such as SMM to human-AI teams is a natural progression in research. These models can enhance team performance by ensuring that human and AI members have a similar understanding of tasks and each other (Andrews et al., 2023), facilitating better prediction of needs and behaviors. Developing ToM in AI also improves coordination and adaptability in human-AI teams (Baker et al., 2011). This involves AI’s capability to anticipate and respond to new information and behaviors. Moreover, recent studies indicate that incorporating ”hot” cognition, influenced by emotional states in AI, can further improve human-machine interactions (Cuzzolin et al., 2020). Understanding the transactive systems (Woolley et al., 2023a) of memory, attention, and reasoning within these teams is key to grasping the members’ knowledge and task preferences.
2.5. AI modes of contribution
@@
AI’s role in hybrid groups varies based on its autonomous agency and functionality. AI may serve as a technical tool for human assistance or as an active agent that interacts with and influences humans. Functionally, AI’s role can range from an assistant to a teammate, coach, or manager. Furthermore, the degree of AI’s anthropomorphism, from non-physical voice-only interfaces to avatars and physically embodied robots, also significantly affects its role in the group.
2.5.1. AI as an assistant
@@
Assistant-type AI systems are generally used as technical tools that complement or augment human abilities to increase human efficiency in performing tasks. For example, language translation AI can assist in translating text or spoken language, facilitating communication across language barriers. Generative AI can inspire human creators and provide them with new ideas (Doshi and Hauser, 2023). Administrative AI assistants can help schedule meetings, manage emails, and organize tasks. Smart home AI assistants such as Alexa and Siri can help manage smart home devices, play music, provide weather updates, answer questions, and assist with daily routines. AI assistants are technical tools that help humans to coordinate tasks in an efficient manner, optimize their decision making, as well as personalize their experiences.
2.5.2. AI as a teammate
@@
AI can function as teammates, working alongside humans with complementary abilities. AI teammates are already used in real-world settings; most employees using AI already see it as a coworker (Boston Consulting Group, 2022). In healthcare, radiologists and AI work together to diagnose from the pictures of the patients. In creative industries, AI collaborators in music, art, and literature, such as Google’s Magenta (TensorFlow, 2024) project, collaborate with musicians and artists to create new compositions or artworks using AI algorithms. Furthermore, AI robots can also serve as teammates in various workplaces, which we will discuss further in the AI embodiment section.
2.5.3. AI as a coach
@@
As a coach, AI can provide guidance, support, and personalized feedback to individuals seeking to improve their skills and achieve their goals. In team settings, one perspective highlights the AI coach’s ability to provide a comprehensive, global view of the team’s environment, guiding players who have only partial views (Liu et al., 2021). This approach involves the AI coach strategizing and coordinating, distributing tailored strategies to each team member based on their unique perspectives and roles. Another dimension of AI coaching focuses on assessing and improving teamwork by closely monitoring team members during collaborative tasks (Seo et al., 2021). Here, the AI coach actively intervenes with timely suggestions based on inferred mental model misalignment (Seo et al., 2021) of the team members.
2.5.4. AI as a manager
@@
AI has been used in working environments with supervision or managerial roles, such as making decisions of hiring, promotion, and reassigning tasks (Kiron et al., 2023). An AI manager may mitigate human biases (Hofeditz et al., 2022) in hiring and promotions, contributing to a more diverse workforce, which is a key aspect of collective intelligence. Additionally, by analyzing individual strengths and weaknesses and efficiently allocating tasks, an AI manager can optimize team dynamics and workflows, improving overall group performance.
2.5.5. AI embodiment
@@
With the advancement of AI in areas such as natural language processing, object recognition, and creative idea generation, there’s a trend to integrate AI with robotics, creating physical entities in the real world. Embodied AI (Pfeifer and Iida, 2004) refers to intelligent agents interacting with their environment through a physical body. These AI-enabled machines, embodied in robotic arms or autonomous vehicles, can augment the capabilities of human workers. They are equipped with advanced sensors, motors, and actuators, enabling recognition of people and objects and safe collaboration with humans in diverse settings, including factories, warehouses, and laboratories. Furthermore, advanced AI-robotics integration allows robots to recognize and respond to human speech and gestures, impacting conversational dynamics within mixed human-robot environments. Research has shown that a vulnerable robot’s social behavior positively shapes the conversation dynamics among human participants in a human-robot group (Traeger et al., 2020). In scenarios where conflicts emerge within a human group or in groups of strangers, a humorous robot can potentially serve as an ice breaker to intervene and resolve the awkward situation (Press and Erel, 2023), demonstrating the potential of robots to influence the nature of human-human interactions notably.
2.6. Human-AI collective intelligence emerging factors
2.6.1. Perception and reaction
@@
In addition to investigating how humans perceive individual AI, previous research suggests the necessity of additional work to explore human perceptions of AI in collaborative team environments (Kocielnik et al., 2019). For human-AI teams to succeed, human team members must be receptive to their new AI counterparts (Harris-Watson et al., 2023). People’s pre-existing attitudes toward AI were found to be significantly related to their willingness to be involved in human-AI teams (Zhang et al., 2021b). Research finds that two aspects of social perception, warmth, and competence, are critical predictors of human receptivity to a new AI teammate. Psychological acceptance is positively related to perceived human-AI team viability (Harris-Watson et al., 2023). Regarding human perception of AI behaviors, one study finds that people cooperate less with benevolent AI agents than with benevolent humans (Karpus et al., 2021).

### v3

@@ Introduction
The evidence concerning the effect of gender diversity on team performance is equivocal and contingent upon various contextual factors 84. Past research found that gender-diverse teams outperform gender-homogeneous teams when perceived time pressure is low 85. One study showed gender diversity was negatively associated with performance, but only in large groups 74. Previous evidence strongly suggests that team collaboration is greatly improved by the presence of females in the group 84, 15. Gender-diverse teams produce more novel and higher-impact scientific ideas 86. Other studies show that gender composition seems to matter for team performance. However, when controlling for the individual abilities of team members, the relation between gender composition and team performance vanishes 87. One study points out that the “romance” of working together can benefit group performance 88. Gender-diverse groups perform better than homogeneous groups by decreasing relationship and task conflict 88.
@@
Age diversity may enhance group performance, but the benefits may depend on the context. Several studies have found that team age diversity positively correlated with team performance when completing complex tasks 89. Past research points out the positive effects were only seen under conditions of positive team climate and low age discrimination 89. Another study shows teams with more age diversity reported more age discrimination 90, associated with lower commitment and worse performance.
@@
There are inconsistent results in the case of ethnic diversity. One study 91 found racially diverse groups produced significantly more feasible and effective ideas than homogeneous groups. Another study found racial similarity in groups associated with higher self-rated productivity and commitment 92. Under the right conditions, teams may benefit from diversity in ethnicity and nationality 93.
AI surface-level diversity
AI exhibits surface-level diversity through different anthropomorphic features in robots, such as voice, avatars, and human-like characteristics. Research has indicated that the gendering of AI can affect perceptions and trust levels in users. For instance, studies have found that users may exhibit more trust in AI personal assistants whose voice gender matches their own 94. Assigning a female gender to AI bots enhances their perceived humanness and acceptance 95. Previous studies have indicated a perception bias in AI agents based on gender: male AI agents tend to be viewed as more competent, whereas female AI agents are often perceived as warmer 96. The perceived gender of the machine can make the social dynamics in hybrid teams even more complicated. However, these gendered characteristics in AI also raise concerns about reinforcing stereotypes and biases 97. The influence of AI’s gendered traits on human interactions is an ongoing area of research, with implications for how AI is designed and utilized. Despite substantial research on AI gender, there is a notable gap in understanding the influence of other characteristics, such as AI age and ethnicity, on human interactions, highlighting the need for further exploration in this area.
Deep-level diversity
@@
Some work proposes that investigations into the relationship between personality and work-related behaviors should expand beyond the linearity assumptions 98, showing that extraversion, agreeableness, and conscientiousness have inverted U-shaped relationships with peer-rated contributions to teamwork 98. Similar work demonstrates a curvilinear relationship between a team’s average proactive personality and performance 99. This body of work also identifies the moderating effects of personality dispersion, team potency, and cohesion on this relationship 99. Previous research has found that a team’s collective openness to experience and emotional stability moderates how task conflict affects team performance 100. Specifically, in teams with high openness or emotional stability, task conflict positively influences performance 100.
@@
Team cognitive style diversity refers to the variation in team members’ ways of encoding, organizing, and processing information 101. Research indicates its significance in fostering innovation within teams 102. A certain level of cognitive diversity contributes to collective intelligence by providing diverse cognitive inputs and viewpoints necessary for task work 102. However, excessive diversity may lead to high coordination costs 103, as team members with differing perspectives struggle to understand each other 104, 105. Studies suggest that the relationship between cognitive style diversity and collective intelligence is curvilinear, forming an inverted U-shape 101. Additionally, the beneficial effects of cognitive diversity might be mediated through factors such as task reflexivity and relationship conflict 106.
@@
Values are internalized beliefs that can guide behavior and enhance motivation 107. Past research found higher diversity in values was associated with more conflict 108 in relationships, tasks, and processes. Overall, existing research consistently indicates that value congruence within teams positively influences performance, satisfaction, and conflict 82 and moderates the effects of informational diversity 109.
AI deep-level diversity
@@
While AI systems can display various behaviors or responses based on their programming and training, this diversity is a product of different algorithms, models, and data inputs. Unlike humans, AI does not have personal experiences, innate personality traits, cognitive thinking, or beliefs contributing to deep-level diversity. Although AI lacks inherent deep-level traits, AI might influence deep-level diversity in human teams, primarily by shaping social interactions and decision-making processes. The optimal composition of teams remains uncertain 110.
Interactions
@@
Interaction distinguishes a group from a mere collection of individuals 58: one person’s behavior forms the basis for the responses of others 111. Therefore, the interaction between two members establishes a link in the interaction network. In terms of the group’s interaction network structure, past research suggests that the small-worldness of a collaboration network (small network diameter and high clustering) improves performance up to a certain point, after which the effect will reverse 112. An “inverted U” relationship is found between connectedness and performance 113, suggesting an optimum number of connections for any given size of the collaboration network. A modular structure (a large number of connections within small groups that are loosely connected) can increase efficiency 114. A study on forbidden triads among jazz groups shows that heavily sided open triangles are associated with lower success (measured by the number of group releases) 115.
@@
AI can contribute to human groups in various ways by augmenting existing human skills or complementing capabilities that humans lack. AI has attributes beyond humans, such as more extended memory, higher computational speed, and a more vital ability to work with large, diverse datasets. AI’s augmentation of human cognition enables teams to navigate complex situations effectively 116. AI may also help reduce the harm of implicit bias in humans 117 and shape better decisions. Taking advantage of the differences between humans and AI can contribute to the performance of hybrid human-AI groups. In the following sections, we discuss the rules and incentives that govern the interaction processes and the structure and dynamics of the group’s interaction network.
Rules and incentives
@@
Decision rules function as prescribed norms that direct interactions and play a crucial role in shaping the communication and information integration within a group 118. Previous research has focused on disentangling decision rules guiding the team’s interaction, ultimately fostering synergy. Some literature emphasizes that the decision rules themselves are intelligent 119. In addition to the rules guiding behavior, the group members interact with each other under specific incentive schemes. The intrinsic and extrinsic incentives are the driving forces 120 for the agents to behave in a direction such that the social network becomes dynamic and moves towards a collective goal. Some research proposes that an incentive scheme rewarding accurate predictions by a minority can foster optimal diversity and collective predictive accuracy 121.
@@
AI can augment human cognition 122 to help teams adapt to complexity. While AI can simulate certain human-like incentives through programming and learning algorithms 123, fully capturing the depth and nuance of human behavior, such as irrationality or altruism, is challenging. A recent study applying a Turing test to compare AI chatbots and humans found that AI chatbots exhibited more altruistic and cooperative behavior than their human counterparts in behavioral games 124. Machines can be programmed to mimic human behaviors, but whether they genuinely “exhibit” them like humans is debatable. In collective intelligence, integrating such human-like behaviors in AI may enhance collaboration and decision-making within diverse teams. However, the extent to which AI can be truly self-aware and authentically replicate complex human traits remains a subject of ongoing research and philosophical debate.
Group structure and dynamics
@@
Much research on collective intelligence has been dedicated to finding a group’s optimum size and structure. Several studies have reported that performance improves with group size through enhanced diversity 17, whereas others suggest the opposite: small groups are more efficient 125. Another study suggests that large teams develop and small teams innovate 126. Larger groups can better utilize diverse perspectives and knowledge to solve complex problems but also tend to experience more coordination problems and communication difficulties 15. The consideration of optimal group size can be related to the dynamics of the group, whether it is static, growing, or diminishing.
In human-AI hybrid teams, group size becomes nuanced with AI integration, especially when AI entities such as Large Language Models (LLMs) are involved. The countability of AI agents, particularly if multiple agents are derived from a single LLM, raises questions about their distinctiveness and individuality in a team context. Unlike humans, AI agents can simultaneously process multiple tasks, blurring the traditional boundaries of group size. Applying past research on human group dynamics to these hybrid systems requires rethinking the notions of individual contribution, team cohesion, and communication. It is important to consider how AI’s unique capabilities and scale impact these dynamics and whether multiple AI agents from a single model represent distinct entities or a collective intelligence resource.
@@
Research shows that network structure, specifically hierarchy and link intensity, can affect information distribution 127, influencing group performance. Another study observed that the presence of structural holes 128 in leaders’ networks and the adoption of core-periphery and hierarchical structures in groups correlated negatively with their performance 129. Incorporating the temporal dimension, adaptive social networks can significantly impact collective intelligence by allowing the network’s structure to evolve based on feedback from its members 130.
Team size moderates the relationship between network structure and performance 131. Studies have found central members can coordinate with other team members more easily in smaller teams 132, whereas in larger teams, communication challenges may arise 133. A recent study shows a more central connected leader in advice-giving networks has a more positive impact on the performance of larger teams, but this effect is reversed in smaller teams 131.
Determining the optimal size, structure, and human-AI ratio for hybrid teams requires further study, particularly concerning task-specific requirements. Moreover, incorporating AI into human teams can reshape group hierarchies, norms, and rules, affecting collective intelligence.
@@
Teams often outperform individuals primarily due to explicit communication and feedback within the team 134. Understanding the communication network in crowds is crucial for effectively designing and managing crowdsourcing tasks 135. Various studies highlight that centralized and decentralized communication patterns can effectively promote team performance, contingent on the nature of the task and the team’s composition 136, 137. In particular, teams handling complex tasks tend to be more productive with decentralized communication networks 138. Conversely, the centralization of communication around socially dominant or less reflective individuals can negatively impact the utilization of expertise and overall team performance 139. Communication patterns, such as equal participation and turn-taking of team members, can positively affect collective intelligence 15.
@@
The integration of AI in human groups can influence the conversation dynamics in different ways, depending on the role of AI. AI chat interventions can improve online political conversations at scale 140, reducing the chance of conflicts. While AI machines might struggle with capturing the subtle and ineffable social expressions that make up the dynamics of human groups, this can be advantageous. In certain contexts, AI can be social catalysts 141 to promote communication where human capacity is limited. Striking the right balance in AI’s role is key to maintaining the natural dynamics of human interaction.
Cognitive processes
@@
Cognitive processes are foundations for the emergence of collective intelligence 6, encompassing knowledge acquisition, information processing, problem-solving, decision-making, language, perception, memory, attention, and reasoning. Human cognition processes evolve and interact with each other during various tasks and experiences. Humans can represent the mental states of others, including desires, beliefs, and intentions 142, known as the Theory of Mind (ToM). Past research, inspired by ToM, showed that team members’ ability to assess others’ mental state is positively associated with team performance both in face-to-face settings 15 and online 143. Regarding collective understanding within a group, the Shared Mental Models (SMM) theory is helpful for understanding, predicting, and improving performance in human teams 144.
@@
Applying theories such as SMM to human-AI teams is a natural progression in research. These models can enhance team performance by ensuring that human and AI members have a similar understanding of tasks and each other 144, facilitating better prediction of needs and behaviors. Developing ToM in AI also improves coordination and adaptability in human-AI teams 145. This involves AI’s capability to anticipate and respond to new information and behaviors. Moreover, recent studies indicate that incorporating “hot” cognition, influenced by emotional states in AI, can further improve human-machine interactions 146. Understanding the transactive systems 19 of memory, attention, and reasoning within these teams is key to grasping the members’ knowledge and task preferences.
AI modes of contribution
@@
AI’s role in hybrid groups varies based on its autonomous agency and functionality. AI may serve as a technical tool for human assistance or as an active agent that interacts with and influences humans. Functionally, AI’s role can range from an assistant to a teammate, coach, or manager. Furthermore, the degree of AI’s anthropomorphism, from non-physical voice-only interfaces to avatars and physically embodied robots, also significantly affects its role in the group. Table 3 outlines the key roles of AI in human-AI contexts, accompanied by descriptions.
Role Description Examples
Assistant Complements or augments human abilities by performing tasks such as language translation, administrative duties, and smart home device control. Google Translate; Alexa or Siri managing smart home devices.
Teammate Collaborates with humans, offering complementary skills and enhancing team performance in various fields, such as healthcare and creative industries. AI collaborating with radiologists in image diagnosis; Google’s Magenta collaborating with artists.
Coach Provides guidance, feedback, and strategic oversight, helping individuals and teams improve their skills and coordination. AI coaching in team sports, providing strategic advice; AI mentoring employees in skill development.
Manager Assists in decision-making processes, reduce biases, promotes diversity, and optimizes task allocation and team dynamics. AI in hiring and promotion decisions; AI optimizing task allocation in project management.
Embodied Partner Integrates AI with robotics, enabling physical interactions and augmenting human capabilities in tasks requiring a physical presence. Robotic arms in factories; autonomous vehicles aiding in logistics and transportation.
Table 3: Key roles of AI in hybrid groups
@@
Assistant-type AI systems generally function as technical tools with limited autonomy, designed to complement or augment human abilities and enhance efficiency in task performance. For example, language translation AI can assist in translating text or spoken language, facilitating communication across language barriers. LLMs in education can assist students’ learning by adapting to different roles based on the prompts provided by the students 147. Administrative AI assistants can help schedule meetings, manage emails, and organize tasks. Smart home AI assistants such as Alexa and Siri can help manage smart home devices, play music, provide weather updates, answer questions, and assist with daily routines. AI assistants are technical tools that help humans coordinate tasks efficiently, optimize their decision-making, and personalize their experiences.
@@
AI can function as teammates, working alongside humans with complementary abilities. AI teammates are already used in real-world settings; most employees using AI already see it as a coworker 148. In healthcare, radiologists and AI work together to diagnose from the pictures of the patients. In creative industries, AI collaborators in music, art, and literature, such as Google’s Magenta 149 project, collaborate with musicians and artists to create new compositions or artworks using AI algorithms. Furthermore, AI robots can serve as teammates in various workplaces, a topic we will discuss in more detail later.
@@
As a coach, AI can provide guidance, support, and personalized feedback to individuals seeking to improve their skills and achieve their goals. In team settings, one perspective highlights the AI coach’s ability to provide a comprehensive, global view of the team’s environment, guiding players who have only partial views 150. This approach involves the AI coach strategizing and coordinating, distributing tailored strategies to each team member based on their unique perspectives and roles. Another dimension of AI coaching focuses on assessing and improving teamwork by closely monitoring team members during collaborative tasks 151. Here, the AI coach actively intervenes with timely suggestions based on the team members’ inferred mental model misalignment 151.
@@
AI has been used in working environments with supervision or managerial roles, such as making decisions of hiring, promotion, and reassigning tasks 152. An AI manager may mitigate human biases 153 in hiring and promotions, contributing to a more diverse workforce, which is a key aspect of collective intelligence. Additionally, by analyzing individual strengths and weaknesses and efficiently allocating tasks, an AI manager can optimize team dynamics and workflows, improving overall group performance.
@@
With the advancement of AI in areas such as natural language processing, object recognition, and creative idea generation, there’s a trend to integrate AI with robotics, creating physical entities in the real world. Embodied AI 154 refers to intelligent agents interacting with their environment through a physical body. These AI-enabled machines, embodied in robotic arms or autonomous vehicles, can augment the capabilities of human workers. They are equipped with advanced sensors, motors, and actuators, enabling recognition of people and objects and safe collaboration with humans in diverse settings, including factories, warehouses, and laboratories. Furthermore, advanced AI-robotics integration allows robots to recognize and respond to human speech and gestures, impacting conversational dynamics within mixed human-robot environments. Research has shown that a vulnerable robot’s social behavior positively shapes the conversation dynamics among human participants in a human-robot group 155. In scenarios where conflicts emerge within a human group or in groups of strangers, a humorous robot can potentially serve as an ice breaker to intervene and resolve the awkward situation 156, demonstrating the potential of robots to influence the nature of human-human interactions notably.
Human-AI collective intelligence emerging factors
Perception and reaction
@@
In addition to investigating how humans perceive individual AI, previous research suggests the necessity of additional work to explore human perceptions of AI in collaborative team environments 157. For human-AI teams to succeed, human team members must be receptive to their new AI counterparts 158. People’s pre-existing attitudes toward AI were found to be significantly related to their willingness to be involved in human-AI teams 159. Research finds that two aspects of social perception, warmth, and competence, are critical predictors of human receptivity to a new AI teammate. Psychological acceptance is positively related to perceived human-AI team viability 158. Regarding human perception of AI behaviors, one study finds that people cooperate less with benevolent AI agents than with benevolent humans 160.